(4.1) Determine the complex numbers i2666 and i145.

(4.2) Let z1 = (6) −i −1+i and z2 = 1+i 1−i . Express z1z3/z2 , z1z2/z3 , and z1/z3z2 in both polar and standard forms.

(4.3) Additional Exercises for practice: Express z1 = −i, z2 = −1 − i √ 3, and z3= − √ 3 + i in polar form and use your results to find z43 /z21 z-12. Find the roots of the polynomials below.

(a) P(z) = z2 + a for a > 0

(b) P(z) = z3 − z2 + z − 1.

(c) Find the roots of z (4) 3 − 1

(d) Find in standard forms, the cube roots of 8 − 8i

(e) Let w = 1 + i. Solve for the complex number z from the equation z4 = w3 . (4.4) Find the value(s) for λ so that α = i is a root of P(z) = z2 + λz − 6.


1
Expert's answer
2021-07-08T09:55:37-0400

(4.1)


"i^{2666}=(i^4)^{666}i^2=-1"

"i^{145}=(i^4)^{36}i=i"

(4.2)


"z_1=\dfrac{-i}{-1+i}=e^{i{3\pi \over 2}}(\dfrac{\sqrt{2}}{2}e^{-i{3\pi \over 4}})"

"=\dfrac{\sqrt{2}}{2}e^{i{3\pi \over 4}}=\dfrac{\sqrt{2}}{2}(\cos\dfrac{3\pi}{4}+i\sin\dfrac{3\pi}{4})=-\dfrac{1}{2}+i(\dfrac{1}{2})"


"z_2=\dfrac{1+i}{1-i}=\sqrt{2}e^{i{\pi \over 4}}(\dfrac{\sqrt{2}}{2}e^{i{\pi \over 4}})"

"=e^{i{\pi \over 2}}=\cos\dfrac{\pi}{2}+i\sin\dfrac{\pi}{2}=i"

"\dfrac{z_1}{z_2}=\dfrac{\dfrac{\sqrt{2}}{2}e^{i{3\pi \over 4}}}{e^{i{\pi \over 2}}}=\dfrac{\sqrt{2}}{2}e^{i{\pi \over 4}}"

"=\dfrac{\sqrt{2}}{2}(\cos\dfrac{\pi}{4}+i\sin\dfrac{\pi}{4})=\dfrac{1}{2}+i(\dfrac{1}{2})"

"\dfrac{z_2}{z_1}=\dfrac{e^{i{\pi \over 2}}}{\dfrac{\sqrt{2}}{2}e^{i{3\pi \over 4}}}=\sqrt{2}e^{-i{\pi \over 4}}"

"=\sqrt{2}(\cos(-\dfrac{\pi}{4})+i\sin(-\dfrac{\pi}{4}))=1-i"

(4.3)


"z_1=-i=\cos(-\dfrac{\pi}{2})+i\sin(-\dfrac{\pi}{2})"

"z_2=-1-i\sqrt{3}=2(\cos(-\dfrac{2\pi}{3})+i\sin(-\dfrac{2\pi}{3}))"

"z_3=-\sqrt{3}+i=2(\cos(\dfrac{5\pi}{6})+i\sin(\dfrac{5\pi}{6}))"

"(z_3)^4=16(\cos(\dfrac{10\pi}{3})+i\sin(\dfrac{10\pi}{3}))"

"(z_1)^2=\cos(-\pi)+i\sin(-\pi))"

"(z_2)^{-1}=\dfrac{1}{2}(\cos(\dfrac{2\pi}{3})+i\sin(\dfrac{2\pi}{3}))"

"\dfrac{(z_3)^4}{(z_1)^2}\cdot(z_2)^{-1}=8(\cos(5\pi)+i\sin(5\pi))=-8"

(a) "P(z)=z^2+a, a>0"


"z^2+a=0=>z_1=-i\sqrt{a}, z_2=i\sqrt{a}"

(b) "P(z)=z^3-z^2+z-1"


"z^3-z^2+z-1=0"

"z^2(z-1)+(z-1)=0"

"z_1=1, z_2=-i, z_3=i"

(c) "z^3-1=0"


"(z-1)(z^2+z+1)=0"

"z_1=1, z_{2,3}=\dfrac{-1\pm i\sqrt{3}}{2}"

"z_1=1, z_2=-\dfrac{1}{2}-i\dfrac{\sqrt{3}}{2}, z_3=-\dfrac{1}{2}+i\dfrac{\sqrt{3}}{2}"

(d)


"8-8i=8\sqrt{2}(\cos(-\dfrac{\pi}{4})+i\sin(-\dfrac{\pi}{4}))"

"k=0: 2^{7/6}(\cos(-\dfrac{\pi}{12})+i\sin(-\dfrac{\pi}{12}))"

"k=1: 2^{7/6}(\cos(\dfrac{7\pi}{12})+i\sin(\dfrac{7\pi}{12}))"

"k=2: 2^{7/6}(\cos(\dfrac{5\pi}{4})+i\sin(\dfrac{5\pi}{4}))=-2^{2/3}-i(2^{2/3})"

(e)


"w=1+i=\sqrt{2}(\cos(\dfrac{\pi}{4})+i\sin(\dfrac{\pi}{4}))"

"w^3=2^{3/2}(\cos(\dfrac{3\pi}{4})+i\sin(\dfrac{3\pi}{4}))"

"z^4=w^3"


"k=0: 2^{3/8}(\cos(\dfrac{3\pi}{16})+i\sin(\dfrac{3\pi}{16}))"

"k=1: 2^{3/8}(\cos(\dfrac{11\pi}{16})+i\sin(\dfrac{11\pi}{16}))"

"k=2: 2^{3/8}(\cos(\dfrac{19\pi}{16})+i\sin(\dfrac{19\pi}{16}))"

"k=3: 2^{3/8}(\cos(\dfrac{27\pi}{16})+i\sin(\dfrac{27\pi}{16}))"

(4.4)


"P(z)=z^2+\lambda z-6"

"z=i"


"(i)^2+\lambda i-6=0"

"\lambda=-7i"